OpenStudy (anonymous):

L(x) is a linear function. Find L(x), if L(5) = 6 and L(10) = 12

OpenStudy (vishweshshrimali5):

If L(x) is a linear function, then, let L(x) = ax + b. Now using L(5) and L(10) values find out a and b and you will get your answer.

OpenStudy (anonymous):

can you show me an example because im sort of lost

OpenStudy (vishweshshrimali5):

Just put x = 5 in L(x) : L(5) = a(5) + b And then put x = 10 in L(x) L(10) = a(10) + b Subtract : L(10) - L(5) = 10a + b - (5a + b) = \(\left(10a - 5a\right) + \color{blue}{\cancel{b}} - \color{blue}{\cancel{b}}\) = 5a And since, it is given to you that L(10) = 12 and L(5) = 6 Therefore, L(10) - L(5) = 6 But from above work, we got L(10) - L(5) = 5a So, 5a = 6. And a = ?

OpenStudy (vishweshshrimali5):

Can you solve it now ?

OpenStudy (anonymous):

yeah thank you